# A Data-driven Typology of Vision Models from Integrated Representational Metrics

### v2

@@ 4 Discussion
In this work, we systematically compared multiple representational properties across vision models—including geometric organization, tuning of individual neurons, and linearly accessible features—to identify which aspects are universally shared versus model-family specific. Our analysis reveals that geometry-preserving or tuning-preserving metrics (RSA, Soft Matching) strongly discriminate between model families, while Linear Predictivity and CCA-variants shows weaker separation. By applying Similarity Network Fusion from multi-omics analysis, we integrated these diverse metrics, achieving superior separation of model families compared to any single metric. Moreover, hierarchical clustering on the SNF-integrated similarity matrix revealed a data-driven typology of vision models that transcends traditional architectural categories. Just as biological typologies require multiple markers, understanding model representations benefits from integrating complementary perspectives rather than relying on single metrics that capture only partial aspects of representational structure.
We note that SNF is not intended to replace individual similarity metrics, but to provide a stable consensus structure when those metrics disagree. The fused graph captures global, multi-metric regularities in representational organization which could be useful for identifying broad model families, or stable clusters. For mechanistic or fine-grained analysis, researchers can and should examine the individual metrics contributing to the SNF fusion to understand which invariances (e.g., geometric, tuning-curve, subspace) drive a specific similarity. In this way, SNF acts as a robust starting point, with the individual metrics supplying interpretability.
Our analysis has several important limitations. First, we focus primarily on natural image databases (ImageNet, Ecoset, CIFAR) and testing the stability of the discovered typology to more out-of-distribution domains would test the generalizability of these groupings. Second, we focus largely on penultimate layer representations. Though we observe reasonable consistency across layers for most metrics as shown in Figure H.1, comprehensive multi-layer analysis could reveal how representational strategies correspond across network depth. Third, a limitation is that all models used in this study were trained on ImageNet-1k. Exploring how the typology changes for web-scale SSL or image–text models would be an important direction for future work. Finally, while SNF provides a principled integration framework, our typology inherently depends on the choice of input metrics. Future work should systematically evaluate how metric selection influences the discovered groupings and explore alternative integration approaches to validate the robustness of our findings.
This typology framework opens several research avenues. Longitudinal analysis could track how models move through representational space during training, potentially revealing whether all models traverse similar developmental trajectories. Extending the framework to multi-modal models could test whether vision-language models form distinct clusters or integrate into existing groupings. A particularly important direction is comparing our representation-based typology with behavioral groupings. Our research focuses on representational similarity, whereas many “practical implications” of such analyses are actually closer to behavioral/functional similarity, which is another type of model similarity metric. The relationship between behavioral similarity and representational similarity requires further investigation and has been proposed as a complementary desideratum for evaluating representational similarity metrics (Ding et al., 2021; Bo et al., 2024). Do models that cluster together based on internal representations also exhibit similar patterns of errors, biases, or generalization behaviors? If behavioral measures yield different groupings than our SNF-based approach, this would reveal that representational similarity doesn’t necessarily imply functional similarity. Finally, validating whether models within the same representational cluster show similar transfer learning performance (e.g., similar fine-tuning convergence rates or final accuracies on downstream tasks) could provide practical utility for the typology and guide model selection strategies. Our framework provides a tool for the community to assess whether new models offer genuinely novel representational strategies or represent variations on established themes.
@@ Appendix A Experiment Settings
Datasets.
The chosen datasets are balanced across different classes as shown in Tabel A.1.
@@
Models.
All models are trained on the ImageNet-1K training set. We obtain pretrained weights from torchvision (maintainers & contributors, 2016), Torch Hub (Paszke et al., 2019), timm (Wightman, 2019), or the official repositories. Unless noted otherwise, we extract activations from each model’s penultimate layer. For CNNs, which commonly include global average pooling, we use that pooled feature. For ViT-style models, we average non-CLS token embeddings to form the final representation for consistency across architectures.
@@ Appendix B Model Family and Architecture Choices
Convolutional Neural Network (supervised; CNN (sup.)).
Bottom-up hierarchies with convolutions and pooling that impose strong local inductive biases. We include AlexNet (Krizhevsky et al., ), VGG-11/13/16/19 (with/without batch normalization) (Simonyan & Zisserman, ), and ResNet-18/34/50/101/152 (He et al., ).
Transformer (supervised; Trans (sup.)).
Vision Transformers partition images into fixed-size patches and use multi-head self-attention for global interactions (Dosovitskiy et al., 2021). We include ViT-S/16, ViT-B/16, ViT-L/16, and ViT-B/32.
ConvNeXt (Liu et al., 2022).
A convolutional family inspired by Transformer design (e.g., large-kernel depthwise convolutions, patchified stems, inverted bottlenecks). We use ConvNeXt-Tiny/Small/Base/Large.
Swin Transformer (Liu et al., 2021).
A hierarchical Transformer with shifted window attention for efficient locality while retaining global context. We use Swin-Tiny/Small/Base/Large.
Convolutional Neural Network (self-supervised; CNN (selfsup.)).
Methods trained without labels using CNN backbones (ResNet-50). We include MoCo (Chen* et al., 2021), DINO (Caron et al., 2021), SwAV (Caron et al., 2020), and Barlow Twins (Zbontar et al., 2021), spanning contrastive and non-contrastive paradigms (momentum contrast, self-distillation, online clustering, and redundancy reduction).
Transformer (self-supervised; Trans (selfsup.)).
Label-free training with Transformer backbones. We include DINO-ViT-Small/16 and DINO-ViT-Base/16  (Caron et al., 2021), MoCo-ViT-Base/16 (Chen* et al., 2021), and MAE-ViT-Base/16 and MAE-ViT-Large/16 (He et al., 2021).
Appendix C Fine-grained Separation Performance on Other Datasets
@@
Appendix D Mean Separability on Other Datasets
Figure D.1: Same as Figure 3, but using Ecoset instead of ImageNet.
Figure D.2: Same as Figure 3, but using CIFAR10 instead of ImageNet.
Figure D.3: Same as Figure 3, but using CIFAR100 instead of ImageNet.
Appendix E Metrics’ Similarity Scores Consistency
Figure E.1: Each subplot (ImageNet, Ecoset, CIFAR10, CIFAR100) shows pairwise Pearson correlations between the vectorized upper-triangle entries of the symmetrized model–model similarity matrices produced by nine metrics. Higher values indicate that two metrics have more similar relational geometry among models.
Appendix F More Clustering Performance
Figure F.1: Same as Figure 4, but for the other 4 metrics.
Figure F.2: Same as Figure 4 and Figure F.1, but using Ecoset instead of ImageNet.
Figure F.3: Same as Figure 5, but using Ecoset instead of ImageNet.
Figure F.4: Figure 4 and Figure F.1, but using CIFAR10 instead of ImageNet.
Figure F.5: Compared to ImageNet in Figure 5, the clustering result is a little different but similar, demonstrating the SNF results and metric’s separability are also influenced by datasets, while also preserving a certain degree of stability. The supervised models are clustered in the same way, but the clusters for the self-supervised models changed, demonstrating that the self-supervised way leads to a special but kind of unifying representational geometry.
Figure F.6: Figure 4 and Figure F.1, but using CIFAR100 instead of ImageNet.
Figure F.7: Same as Figure F.5, but using CIFAR100 instead of CIFAR10.
Appendix G Cophenetic Correlation Coefficients for Clustering
Figure G.1: Cophenetic correlation coefficients (CCC) for hierarchical clusterings induced by each metric on four datasets. Higher CCC (closer to 11) means the clustering more faithfully preserves the original pairwise structure. Bars show CCC on ImageNet, Ecoset, CIFAR10, and CIFAR100 (values annotated; lines trace dataset-wise trends).
Appendix H Metrics’ Cross Layer Consistency
@@
Figure H.1: Cross-layer consistency of inter-model similarity. The bar height is the mean of the three depth pairs correlations; higher values indicate greater consistency.
Appendix I Brain Clustering Validation
Figure I.1: Cross-regional relationships were derived from five similarity measures using PCA analysis on NSD data. Each point represents a brain region instance, and text labels indicate centroid positions. SNF fusion shows the best intra-class compactness and inter-class separation.
A.
B.
Figure I.2: A. Mean $d^{\prime}$ score on NSD. B. Significant hierarchical scores comparison of different representations across the ventral visual pathway (NSD, $n=1000$, $p<0.05$).
For brain data, we used responses to the 1,000 shared natural images from the Natural Scenes Dataset (NSD) (Allen et al., ), spanning 10 visual regions across 4 subjects. We exclude SVCCA and PWCCA from brain analyses because their alignment values were inflated by the high voxel count relative to the number of stimuli.
To assess brain region discriminability, we quantified the difference between within-region and across-region representational similarities using the separability measure $d^{\prime}$. SNF which integrates information across all representational dimensions, achieves dramatically superior family separation compared to any single metric. In Figure I.2A, SNF attains a mean $d^{\prime}$ of 21.45—nearly five times higher than the best-performing single measure—and consistently outperforms all baselines across separation criteria. To visualize and corroborate these patterns, we applied PCA to each similarity matrix. All metrics recovered broad trends (Figure I.1), but the SNF-fused matrix produced the clearest, most interpretable layout. Early visual areas (V1-V4) traced a smooth trajectory, with V4 in a transitional position between the early cortex and higher-level ventral, dorsal, and lateral streams.
To quantify how well a representation preserves the expected ventral-stream hierarchy, as suggested in Thobani et al. (2025), we use a hierarchy score based on Kendall’s $\tau$. Each region $r$ is assigned a discrete hierarchical level $L(r)\in\{1,\dots,5\}$, as $L(\text{V1v})=L(\text{V1d})=1,L(\text{V2v})=L(\text{V2d})=2,L(\text{V3v})=L(\text{V3d})=3,L(\text{V4})=4,L(\text{ventral\_visual})=5$, given a brain-region-by-region similarity matrix $S\in\mathbb{R}^{R\times R}$ and the corresponding region labels $\{r_{1},\dots,r_{R}\}$, we first construct a hierarchy-distance matrix,
$H_{ij}=\bigl|L(r_{i})-L(r_{j})\bigr|,\quad 1\leq i,j\leq R.$
We then extract all off-diagonal entries of $H$ and $S$,
$\mathbf{h}=\{H_{ij}:i\neq j\},\qquad\mathbf{s}=\{S_{ij}:i\neq j\},$
and define $\mathbf{y}=-\mathbf{s}$ so that smaller hierarchy distance corresponds to larger similarity. The hierarchy score is the Kendall’s $\tau$-b rank correlation between $\mathbf{h}$ and $\mathbf{y}$: $\tau_{\text{hier}}=\tau_{\text{Kendall}}\!\bigl(\mathbf{h},\mathbf{y}\bigr)$. These results further support the validity of SNF in revealing known biological correspondence structure (Figure I.2B).
Appendix J The Usage of Large Language Models (LLMs)

### v4

@@ 4 Discussion
This work reframes representational comparison as a question of what kinds of structure different similarity metrics are capable of recovering. Rather than treating representational similarity as a single quantity, we systematically evaluated multiple correspondence dimensions (e.g. geometric organization, unit-level tuning, and linearly accessible information) and assessed how well each recovers known structure in two domains: procedural differences among artificial vision models and anatomical–functional organization in human visual cortex. Across both domains, we find a consistent ordering of representational similarity measures. Metrics that preserve representational geometry or unit-level tuning (e.g., RSA, linear CKA, SoftMatch) most reliably recover known structure, distinguishing procedurally distinct model families and differentiating cortical regions across subjects. In contrast, more flexible mappings that emphasize linearly accessible information (e.g., Linear Predictivity and CCA-based methods) show substantially weaker discrimination. These results suggest that while linearly decodable signals may reflect more broadly shared constraints, finer-grained geometric and tuning structure carries stronger signatures of procedural and anatomical organization.
Another contribution of this work is to show that these representational dimensions are complementary rather than redundant. No single metric provides a complete characterization of representational correspondence. To address this, we adapt Similarity Network Fusion (SNF), originally developed for multi-omics integration, to combine similarity graphs derived from multiple metrics. The fused similarity consistently outperforms individual measures in recovering known structure in both models and brains. We note that SNF provides a stable global summary when different metrics capture complementary aspects of representation, but does not replace individual similarity measures. While the fused similarity is well-suited for identifying broad organizational structure, such as model families or cortical hierarchies, individual metrics are necessary to interpret which representational properties drive the observed similarity. More broadly, our results suggest that representational similarity measures should be interpreted in terms of the structure they are capable of recovering, rather than interchangeable proxies for representational correspondence.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Experiment Settings
Datasets. The chosen datasets are balanced across different classes as shown in Tabel A.1.
@@
Models. All models are trained on the ImageNet-1K training set. We obtain pretrained weights from torchvision (maintainers & contributors, 2016), Torch Hub (Paszke et al., 2019), timm (Wightman, ), or the official repositories. Unless noted otherwise, we extract activations from each model’s penultimate layer. For CNNs, which commonly include global average pooling, we use that pooled feature. For ViT-style models, we average non-CLS token embeddings to form the final representation for consistency across architectures.
@@ Appendix B Model Family and Architecture Choices
Convolutional Neural Network (supervised; CNN (sup.)). Bottom-up hierarchies with convolutions and pooling that impose strong local inductive biases. We include AlexNet (Krizhevsky et al., ), VGG-11/13/16/19 (with/without batch normalization) (Simonyan & Zisserman, ), and ResNet-18/34/50/101/152 (He et al., ).
Transformer (supervised; Trans (sup.)). Vision Transformers partition images into fixed-size patches and use multi-head self-attention for global interactions (Dosovitskiy et al., 2021). We include ViT-S/16, ViT-B/16, ViT-L/16, and ViT-B/32.
ConvNeXt (Liu et al., 2022). A convolutional family inspired by Transformer design (e.g., large-kernel depthwise convolutions, patchified stems, inverted bottlenecks). We use ConvNeXt-Tiny/Small/Base/Large.
Swin Transformer (Liu et al., 2021). A hierarchical Transformer with shifted window attention for efficient locality while retaining global context. We use Swin-Tiny/Small/Base/Large.
Convolutional Neural Network (self-supervised; CNN (selfsup.)). Methods trained without labels using CNN backbones (ResNet-50). We include MoCo (Chen* et al., 2021), DINO (Caron et al., 2021), SwAV (Caron et al., 2020), and Barlow Twins (Zbontar et al., 2021), spanning contrastive and non-contrastive paradigms (momentum contrast, self-distillation, online clustering, and redundancy reduction).
Transformer (self-supervised; Trans (selfsup.)). Label-free training with Transformer backbones. We include DINO-ViT-Small/16 and DINO-ViT-Base/16  (Caron et al., 2021), MoCo-ViT-Base/16 (Chen* et al., 2021), and MAE-ViT-Base/16 and MAE-ViT-Large/16 (He et al., 2021).
Appendix C Separation Performance on Other Datasets
@@
Appendix D Metrics’ Similarity Scores Consistency
Figure D.1: Each subplot (ImageNet, Ecoset, CIFAR10, CIFAR100) shows pairwise Pearson correlations between the vectorized upper-triangle entries of the symmetrized model–model similarity matrices produced by nine metrics. Higher values indicate that two metrics have more similar relational geometry among models.
Appendix E More Clustering Performance
Figure E.1: Hierarchical clustering of models using three functionally distinct representational similarity metrics and a similarity-metric–averaging baseline (see Fig. E.2 for additional metrics). Clustering is performed with average linkage and optimal leaf ordering, based on induced distances (11- similarity score). Rows and columns are deliberately re-ordered to match the leaf ordering produced by the clustering algorithm. Lighter colors indicate higher similarity; diagonal entries (self-comparisons) are omitted.
Figure E.2: Same as Figure E.1, but for the other 4 metrics.
Figure E.3: Same as Figure E.1, but using Ecoset instead of ImageNet.
Figure E.4: Same as Figure E.2, but using Ecoset instead of ImageNet.
Figure E.5: Same as Figure 4, but using Ecoset instead of ImageNet.
Figure E.6: Same as Figure E.1, but using CIFAR10 instead of ImageNet.
Figure E.7: Same as Figure E.2, but using CIFAR10 instead of ImageNet.
Figure E.8: Compared to ImageNet in Figure 4, the clustering result is a little different but similar, demonstrating the SNF results and metric’s separability are also influenced by datasets, while also preserving a certain degree of stability. The supervised models are clustered in the same way, but the clusters for the self-supervised models changed, demonstrating that the self-supervised way leads to a special but kind of unifying representational geometry.
Figure E.9: Same as Figure E.1, but using CIFAR100 instead of ImageNet.
Figure E.10: Same as Figure E.2, but using CIFAR100 instead of ImageNet.
Figure E.11: Same as Figure E.8, but using CIFAR100 instead of CIFAR10.
Appendix F Cophenetic Correlation Coefficients for Clustering
Figure F.1: Cophenetic correlation coefficients (CCC) for hierarchical clusterings induced by each metric on four datasets. Higher CCC (closer to 11) means the clustering more faithfully preserves the original pairwise structure. Bars show CCC on ImageNet, Ecoset, CIFAR10, and CIFAR100 (values annotated; lines trace dataset-wise trends).
Appendix G Metrics’ Cross Layer Consistency
@@
Figure G.1: Cross-layer consistency of inter-model similarity. The bar height is the mean of the three depth pairs correlations; higher values indicate greater consistency.
Appendix H The Usage of Large Language Models (LLMs)